OpenStudy (anonymous):

A(r) is an area function of a circle. Find A(2) when A(r)=3.14r^2

OpenStudy (anonymous):

What is Trial and error method..? Solve (0.36x10^-6) x^3 + 0.45x - 3000 by this method.

OpenStudy (anonymous):

huh?

OpenStudy (jlastino):

replace r with 2 in the equation then calculate :D

OpenStudy (anonymous):

is it 6.28?

OpenStudy (jlastino):

no...the value of r is squared then it is multiplied by 3.14 what is 2 squared?

OpenStudy (anonymous):

4? lol

OpenStudy (jlastino):

yeah then multiply it by 3.14 haha

OpenStudy (anonymous):

oh wait so its A(r)=3.14r^2 A(2)=3.14(2)^2 2^2=4 *3.14? thats how i do this?

OpenStudy (anonymous):

which is is 12.56!!!!!

OpenStudy (anonymous):

omg i got it right!!! lmfao thanks!